40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 14 08:25:29 2014, Steve, Update, SUS, morning issues morning.png
    Reply  Fri Nov 14 11:07:59 2014, manasa, Update, CDS, Not able to run models on FE machines CDSissue.png
       Reply  Fri Nov 14 15:45:34 2014, Jenne, Update, CDS, Computers back up after reboot 
Message ID: 10717     Entry time: Fri Nov 14 15:45:34 2014     In reply to: 10715
Author: Jenne 
Type: Update 
Category: CDS 
Subject: Computers back up after reboot 

[Jenne, Q]

Everything seems to be back up and running.

The computers weren't such a big problem (or at least didn't seem to be).  I turned off the watchdogs, and remotely rebooted all of the computers (except for c1lsc, which Manasa already had gotten working).  After this, I also ssh-ed to c1lsc and restarted all of the models, since half of them froze or something while the other computers were being power cycled. 

However, this power cycling somehow completely screwed up the vertex suspensions.  The MC suspensions were fine, and SRM was fine, but the ITMs, BS and PRM were not damping.  To get them to kind of damp rather than ring up, we had to flip the signs on the pos and pit gains.  Also, we were a little suspicious of potential channel-hopping, since touching one optic was occasionally time-coincident with another optic ringing up.  So, no hard evidence on the channel hopping, but suspicions.

Anyhow, at some point I was concerned about the suspension slow computer, since the watchdogs weren't tripping even though the osem sensor rmses were well over the thresholds, so I keyed that crate.  After this, the watchdogs tripped as expected when we enabled damping but the RMS was higher than the threshold.

I eventually remotely rebooted c1sus again.  This totally fixed everything.  We put all of the local damping gains back to the values that we found them (in particular, undoing our sign flips), and everything seems good again.  I don't know what happened, but we're back online now. 

Q notes that the bounce mode for at least ITMX (haven't checked the others) is rung up.  We should check if it is starting to go down in a few hours.

Also, the FSS slow servo was not running, we restarted it on op340m.

ELOG V3.1.3-